OpenStudy (anonymous):

How far apart are the foci of an ellipse with a major axis of 26 feet and a minor axis of 24 feet?

OpenStudy (anonymous):

a^2+b^2=c^2

OpenStudy (anonymous):

do i just put the numbers into the equation? :/ these are the answer choices:4 feet 5 feet 10 feet 20 feet

OpenStudy (anonymous):

If the major axis is 26, then 2a is 26 and a is 13. If the minor axis is 24, then 2b is 24 and b is 12. Use the values of a and b to calculate c

OpenStudy (anonymous):

do you know the Pythagorean triples or no

OpenStudy (anonymous):

hmm no i don't

OpenStudy (anonymous):

what do you mean 2a? a? 2b?

OpenStudy (anonymous):

The distance between the focal points is 2a=26 and 2b=24

OpenStudy (anonymous):

so plug here a^2+b^2=c^2

OpenStudy (anonymous):

The length of the major axis is 2a and that of the minor axis is 2b like @begzat said. Find the eccentricity of the ellipse sing \[e ^{2}= 1 - b ^{2}/a ^{2}\] [when a>b] .. the distance between the focii is 2ae

OpenStudy (anonymous):

using*

OpenStudy (anonymous):

thanks :)